OBSERVATION AND ANALYSIS OF THE 2ν_{3} BAND OF ^{32}S^{16}O_{3}

\underline{THOMAS A. BLAKE}, ROBERT L. SAMS, Pacific Northwest National Laboratory, P. O. Box 999, Mail Stop K8-88, 3020 Q Avenue, Richland, WA 99352 (PNNL is operated for the US Department of Energy by the Battelle Memorial Institute under contract DE-AC06-76RLO 1830); ARTHUR MAKI, 15012 24th Ave. S. E. Mill Creek, WA 98012; TONY MASIELLO, JEFFREY BARBER, NICO-LAE VULPANOVICI, ENGELENE t.H. CHRYSOSTOM, JOSEPH W. NIBLER, Department of Chemistry, Oregon State University, 153 Gilbert Hall, Corvallis, OR 97331-4003; ALFONS WE-BER, National Science Foundation, 4201 Wilson Boulevard, Arlington, VA 22230.

We are engaged in a comprehensive investigation of the fundamental and low lying combination and overtone bands of the parent ^{32}S^{16}O_{3} as well as the ^{34}S^{16}O_{3}, ^{32}S^{18}O_{3}, and ^{34}S^{18}O_{3} isotopic variants of sulfur trioxide. CARS spectroscopy of the ν_{1} Raman active bands has been performed at Oregon State University and high resolution infrared absorption spectroscopy of room temperature static cell samples has been performed at PNNL using a Bruker IFS 120 HR Fourier transform spectrometer. Results for the absorption spectrum of the 2ν_{3} band of^{32}S^{16}O_{3}, recorded at a resolution of 0.0025 cm^{−1}, will be presented. The overtone of the ν_{3} fundamental [(e′)^{2} = A′_{1} + E′] is centered at 2777.87139(5) cm^{−1}with B′ = 0.3462802(2) cm^{−1}and B′′ = 0.34854336(4) cm^{−1}. Assignment and fitting of the perpendicular band shows that there is an internal perturbation that arises from a Δk = ±2, Δl = ±2 coupling of the l = 0 and l = ±2 components. A level crossing is evident with the l = 2, k = 11 levels pushed to higher energy and the l = 2, k = 14 levels pushed to lower energy, for example. We have determined that the band center of the perturbing dark state is approximately 11 cm^{−1}below that of the bright state. The Hamiltonian including perturbing terms and fit constants will be presented. In addition, ν_{2}, ν_{3}, and ν_{4} have been measured and analyzed to reveal new details such as the splitting of the K = 2 and K = 3 levels.